OpenStudy (anonymous):

My questions are listed below, the questions are there and my attempted answers are in asterix, sorry about the pictures (diagrams) being upside down. I am quite unsure whether these are right if someone could help me that would be great

OpenStudy (anonymous):

I negatively charge a plastic ruler by rubbing it with a piece of material. I have a metal ball hanging from a piece of fishing line (which is an insulator, so charge cannot travel along the fishing line). I move the charged ruler near to the metal ball, without touching it, so that a charge is induced on the metal ball. Does this mean that the net (total) charge on the metal ball is: (a) positive, (b) negative, or (c) remains the same? Explain your answer. ****It would remain the same as neither electrons can leap through air from one object to another*****

OpenStudy (anonymous):

I have two metal balls, A and B, each suspended from a length of fishing line. Both balls have been negatively charged, but ball B has twice the total charge of ball A. When the balls are brought near each other they can be seen to repel each other. How does the force exerted on ball A by ball B compare to that exerted by ball B on ball A? i.e. is it the same, greater, lesser? Explain your answer. ****Fab=-Fba, so equal in magnitude but opposite in direction****

OpenStudy (anonymous):

Draw the electric potential lines (dashed lines) and electric field lines (solid, with arrows) for a parallel plate capacitor; draw at least five of each to make sure you communicate the details of the shape of the potential and the field.

OpenStudy (anonymous):

4.Draw a cicuit diagram of a 6V battery connected across a 10 k? resistor and 1 mF capacitor, in series

OpenStudy (anonymous):

a)When you first connect the battery to the components, what is the potential difference across the resistor? *****I=V/R, I=6/10,000,I=0.0006A=0.6mA. Then Voltage across resistor, alphaV=IxR, alphaV=0.0006Ax10,000=6V, answer=6V*****

OpenStudy (anonymous):

b)Once current has stopped flowing through the circuit, what is the potential difference across the resistor? Explain your answer. *****Potential difference is equal to zero as current equals zero, as alphaV=IxR, when I=0, alphav=0*****